# Consistent Validation for Predictive Methods in Spatial Settings

### v2

@@ E.3.2 Loss Function
which is bounded by $1.0$. The empirical risk is calculated as in Equation 112.
E.3.3 Model Fitting
We fit a Gaussian process regression model to using only the first covariate $\chi^{(1)}$ to the training data. The prior is taken to be the same as the data generating process, but with (only) $\chi^{(1)}$ in place of $(\chi^{(1)},\chi^{(2)})$ in the second kernel in Equation 125 and Equation 127 for the two datasets respectively. The mean of the posterior process is used for predictions, and is calculated using GPFlow [Matthews et al., 2017].
E.3.4 Implementation of Risk Estimation
The holdout is implemented by taking an (unweighted) average of the loss on each validation point. Both nearest neighbor methods are implemented using scikit-learn [Pedregosa et al., 2011] with $kd$-trees and Euclidean distance. For $k^{\star}_{T_{2}}$ nearest neighbors, nearest neighbors is performed for all $k$ that are powers of $2$ less than $N^{\textup{{val}}}$, and the value of $k$ with the smallest bound is used for risk estimation. This is done with $\delta=r$, with $r$ calculated as in Section D.5.4, and $\Delta=1$. Table 2 and Table 3 show the values of $k$ chosen for grid and point prediction respectively. For the grid prediction task, $k^{\star}_{T_{2}}$ tends to be small, even as the size of the validation set becomes large. This supported by Corollary D.16, since even $1$-nearest neighbor reliably estimates risk in this setting. For the point prediction task, $k^{\star}_{T_{2}}$ grows with $N^{\textup{{val}}}$.
@@ E.3.4 Implementation of Risk Estimation
E.3.5 Monte Carlo Estimation of Test Risk
Following the argument in Section E.1, we use the empirical test risk Equation 112 in place of the test risk as ground truth in synthetic experiments. The assumption that $(Z_{m})_{m=1}^{M^{\textup{{test}}}}$ are independent holds by the description of the data generating process, because the $(Y_{m})_{m=1}^{M^{\textup{{test}}}}$ are conditionally independent and $Z_{m}$ is a function of $Y_{m}$. The assumption that $Z_{m}$ is almost surely bounded holds with $\Delta=1$ by our choice of truncated squared loss. Further, in both synthetic experiments, we take $M^{\textup{{test}}}=45000$. Combining these gives that with probability at least $0.95$
@@ E.3.5 Monte Carlo Estimation of Test Risk
Figure 1 shows the absolute difference between each estimator and the empirical test risk across 100 seeds. We see that the difference between the estimators is generally larger than twice Equation 129, and so by the argument in Section E.1, we expect our estimate of ground truth to be accurate enough that the difference in performance of the methods is not simply due to error in estimating the ground truth.
E.3.6 Metrics Reported and Additional Figures
Figures 5 and 7 show the relative errors of each estimation, calculated as
@@ E.3.6 Metrics Reported and Additional Figures
E.4 Air Temperature Tasks
We now provide additional details about data source, pre-processing, model fitting and risk estimation for the air temperature dataset. These are identical between the real response experiment and the partially synthetic experiment, except for the bootstrapping procedure described in Section E.4.6. We also provide additional experimental results on a grid prediction task for both the bootstrapped and original datasets.
E.4.1 Data Sources
The land surface temperature used is from MODIS Aqua [Wan et al., 2021] and is monthly average land surface temperature on a $0.05$ degree grid. We download monthly average weather station data from the Global Historical Climatology Network [Menne et al., 2018]. Latitude and longitude of major United States (US) urban areas are from the 2023 US census gazetteer [United States Census Bureau, 2023]. All of these datasets are produced in large part by US government agencies (NASA, NOAA and the Census). While we could not find specific license information, we understand these datasets to be public domain following section 105 of the Copyright Act of 1976.
E.4.2 Data Pre-processing
We assign the land surface temperature at the nearest point (using a spherical approximation to distance between points) to each weather station. The nearest point is found using the scikit-learn implementation of nearest neighbor algorithm using the ‘ball-tree’ [Omohundro, 2009] and ‘Haversine’ metric. Temperatures are converted to Celsius from Kelvin. We remove all rows where the Land Surface temperature is not available. We use the weather station data uploaded to GHCNM as of January 15, 2024. We filter out weather stations outside of the United States (based on the station ID). We also filter out stations with a non-empty quality control flag or no temperature recorded for January 2018 (the month we consider). Finally, we remove stations in Hawaii or Alaska to focus on the continental United States. In total, after this processing, there are 6422 weather stations. We use 70% of the stations for fitting the models, and holdout the remaining 30% estimates. When building the test sites, we remove points outside the United States based on a reverse geocoding lookup with Thampi [2015] to the nearest city. This does not create an exact boundary (since it is based on the nearest city or town and not the country in which the latitude and longitude is based in) but is a good proxy for whether or not a point is in the United States. Figure 8 of the available weather stations for model fitting and validation, colored by monthly average temperature in January 2018.
@@ E.4.3 Loss Function
E.4.4 Model Fitting
Inspired by Hooker et al. [2018], we fit a geographically weighted least squares regression model using the land surface temperature at day and night. In particular, we fit an affine model, with the coefficients, $\beta(S)$ depending on the location that will be predicted at. $\beta(S)$ is selected by solving the weighted least squares problem,
@@ E.4.4 Model Fitting
E.4.5 Risk Estimation Details
The holdout is implemented as in previous experiments (Section E.3.4). We estimate the standard error of the holdout empirically by computing the sample standard deviation of the sum of the losses,
@@ E.4.6 Bootstrapping of Residuals
We can then directly estimate the test risk via generating many $Y^{\textup{{test}}}$ at each spatial location (we use 1000 realizations for each city in the 5-metros task) and 1 for each grid point in the grid task (since the error is averaged over test sites this still results in an estimator that is concentrated) in this manner and forming a Monte Carlo estimate as in Section E.1.
E.4.7 Estimation of Ground Truth in Bootstrapped Experiment
Following the argument in Section E.1, we use the empirical test risk Equation 112 in place of the test risk as ground truth in synthetic experiments. For the assumption that $(Z_{m})_{m=1}^{M^{\textup{{test}}}}$ are independent to hold it is sufficient for $Y^{\textup{{test}}}_{m}$ to be independent, conditioned on the spatial location at which it is observed. This holds based on the data generating process used to construct the synthetic responses: since $Y^{\textup{{test}}}_{m}$ is a noisy observation of the smooth function we fit to the weather stations, plus noise sampled independently from the distribution of residuals. Because the loss is truncated MAE, the $Z_{m}$ are surely bounded by $5$. We use $10000$ samples at each of the 5 test location in estimating the risk. Using these numbers in Equation 115, we arrive at
@@ E.4.7 Estimation of Ground Truth in Bootstrapped Experiment
Following the argument in Section E.1, we expect our estimate of ground truth to be good enough to distinguish between the quality of models whose absolute error from the estimated ground truth differs by more than $2\times 0.031=0.062$. Figure 9 shows these absolute errors. We see that for many of the seeds, the difference between the performance of 1NN (orange) and SNN and the holdout is greater than $0.174$. Given that the earlier argument is quite conservative (in the sense that Hoeffding’s inequality is likely to be loose), we therefore can attribute the difference in performance of the methods to indicate that SNN and the holdout are giving better estimates of the ground truth test risk, and the observed difference is not due to error in our estimation of the test risk.
@@
We therefore expect our estimate of ground truth to be very accurate, although we see in Figure 10 that the methods all perform well in estimating the test risk on this task, and so there is likely not a meaningful difference in which approach is used to perform validation.
E.4.8 Results for Grid Prediction with Bootstrapped Data
Figure 10 shows the results for holdout, $1$-nearest neighbor and SNN with the test set each grid point in the map that is located in the continental United States. All 3 methods lead to reasonably accurate estimates of the mean absolute error on this prediction task (within 0.1 degrees of the ground truth error). Based on our theory, we generally expect 1NN and SNN to have small error in grid prediction tasks (at least with sufficient data and the infill assumption being satisfied), while for the holdout it depends on the particular predictive method and distribution of test and validation sites. In this case, it appears for both prediction methods the bias introduced by the use of the holdout is relatively small.
@@ E.4.8 Results for Grid Prediction with Bootstrapped Data
E.4.9 Results for Grid Prediction with Real Data
Table 4 shows the results for holdout, $1$-nearest neighbor and SNN with the test set each grid point in the map that is located in the continental United States. We see good agreement between all three method. This is expected for $1$-nearest neighbor and SNN based on earlier theory (Section D.6.5).
@@ E.5 UK Housing Experiment
E.5.1 Data Sources and Pre-processing
We download 2023 price paid data for England and Wales from HM Land Registry [2023]. This data is subject to a UK Open Government License (https://www.nationalarchives.gov.uk/doc/open-government-licence/version/3/), which requires citation of the data, but allows both commercial and non-commercial uses. These records contain postal codes for each property sold, the type of property sold, town or city, price paid for the property. We use the type of property variable to filter out all properties that are not flats, and only consider additions (not replacements or deletions) to the dataset and “standard” price paid data (not repossessions, buy-to-lets or other sales labelled as non-standard). Noting that a postal code in the UK corresponds to a very small geographic area, we obtain latitude and longitude data for each sale by looking up the postal code coordinates using the National Statistics Postcode Lookup [Office Of National Statistics, 2024], which we understand to be a product of the UK Census and therefore also subject to an Open Government License. We convert from northing and easting to latitude and longitude using R. We log transform the price variable prior to model fitting as we expect price paid to be non-negative and highly skewed, and so a Gaussian (process) prior would otherwise be almost certainly inappropriate.
E.5.2 Model Fitting
@@
We fit a Gaussian process regression model with prior covariance specified by a sum of two Matérn 3/2 kernels and a zero prior mean on the mean centered log price paid data. We use a sum of Matérn 3/2 kernel in place of the sum of RBF kernel used in Hensman et al. [2013] as we expect there to be places where (log) property prices vary quickly in space, and so the smoothness properties implicitly assumed in using an RBF kernel may be inappropriate. We use 2000 inducing points for the variational approximation. The locations of these points are optimized jointly with model parameters when maximizing the evidence lower bound. We use the closed form for the optimal variational posterior (given a set of inducing points) provided in Titsias [2009], and perform maximization of the evidence lower bound using L-BFGS.
@@
The locations of the inducing points are initialized by the greedy procedure suggested in Burt et al. [2020], which is essentially equivalent to a partially pivoted Cholesky decomposition recommended earlier in the Gaussian process approximation literature [Foster et al., 2009]. The initial prior variance of both kernels is set to be equal to the variance in the training data; the lengthscales of one kernel (intended to model regional price trends) are initialized to twice the standard deviation in the location data (in radians), while the scale of the other kernel (intended to model local trends) is initialized to half the standard deviation in the location data. The likelihood standard deviation is initialized to be 0.1$\times$ the standard deviation in the log price paid in the training data.
E.5.3 Estimation of Ground Truth
We next describe why we might expect empirical test risk to provide a reasonable estimate of ground truth in this problem, and in particular we discuss the assumptions that justify the use of empirical test risk in relation to Section E.1. We consider a truncated loss $\ell(a,b)=\min(0,10^{6})$, and so $Z_{m}$ is almost surely bounded by $10^{6}$. The assumption that the $(Z_{m})_{m=1}^{M^{\textup{{test}}}}$ are independent would be implied by independence of $(Y_{m})_{m=1}^{M^{\textup{{test}}}}$ that is: that given the location at which a flat is sold, any remaining randomness in the observed sale prices is independent. Concretely, we might think of the randomness in sales price, $\epsilon^{\text{test}}_{m}$ in our model, as coming from aspects of the sale process of the house, such as who happens to see the advertisement for a house, and we assume these are independent for each house when constructing our estimate of the ground truth.
If the conditional independence assumption proposed above holds, then following Section E.1 we have that
@@ E.5.3 Estimation of Ground Truth
E.5.4 Risk Estimation Details
Holdout, 1NN and SNN are run as in the air temperature experiments (Section E.4.5). In particular, we use a failure probability of $\delta=0.1$ for SNN and nearest neighbor calculations are done with respect to Haversine distance to account for Earth’s curvature. We use a fixed $\delta$ since we only discuss a method applicable to estimating fill distance on the unit cube, not on a subset of the sphere. $\Delta=\pounds 1,000,000$ is used in selecting the number of neighbors as this is an upper bound on the truncated loss. We use a Lipschitz constant of £1,000/km as £1/km seems implausibly small.
E.6 Wind Speed Prediction Experiment
@@
E.6.1 Data Sources and Pre-processing
We download daily wind speed readings from weather stations from the Global Historical Climate Network [Menne et al., 2012]. As this dataset was constructed by NOAA employees, we understand it to be public domain following section 105 of the Copyright Act of 1976. We filter out weather stations outside the continental US, as well as any weather stations that do not contain daily average wind speed readings. We look only at wind speed data from January in the prediction task, and years 2000–2024. There is a weather station at Chicago O’Hare which we remove from the training and validation data and use as the test set.
For each replicate used to form Figure 10 we split off a training set containing 80% of weather stations, and a validation set containing the remaining 20%. The number of observations in the training and validation set varies (because different weather stations may be online for a different number of days in January in previous years), but this leads to on the order of 580000 training observation and 126000 validation observations. Each training and validation point is a triple containing latitude, longitude and average wind speed. We perturb the latitude and longitude (in degrees) of validation points by a Gaussian random variable with standard deviation $10^{-12}$, which is essentially equivalent to using random tie-breaking in the nearest neighbor algorithms. We expect this has a significant impact on 1NN compared to the version discussed in the paper, because there may be many observations from the nearest weather station to Chicago O’Hare. While this would unlikely be done in practice, random tie-breaking, or tie-breaking by selecting the first nearest neighbor according to some other ordering are common and would lead to similar outcomes as the results presented here (but higher variance than averaging over all neighbors that are equally close). The latitude and longitude are then converted to radians for the analysis.
E.6.2 Loss Function

### v3

@@ D.3.2 Loss Function
which is bounded by $1.0$. The empirical risk is calculated as in eq. 114.
D.3.3 Model Fitting
We fit a Gaussian process regression model to using only the first covariate $\chi^{(1)}$ to the training data. The prior is taken to be the same as the data generating process, but with (only) $\chi^{(1)}$ in place of $(\chi^{(1)},\chi^{(2)})$ in the second kernel in eq. 127 and eq. 129 for the two datasets respectively. The mean of the posterior process is used for predictions, and is calculated using GPFlow (Matthews et al., 2017).
D.3.4 Implementation of Risk Estimation
The holdout is implemented by taking an (unweighted) average of the loss on each validation point. Both nearest neighbor methods are implemented using scikit-learn (Pedregosa et al., 2011) with $kd$-trees and Euclidean distance. For $k^{\star}_{T_{2}}$ nearest neighbors, nearest neighbors is performed for all $k$ that are powers of $2$ less than $N^{\textup{{val}}}$, and the value of $k$ with the smallest bound is used for risk estimation. This is done with $\delta=r$, with $r$ calculated as in section C.5.4, and $\Delta=1$. Table 2 and table 3 show the values of $k$ chosen for grid and point prediction respectively. For the grid prediction task, $k^{\star}_{T_{2}}$ tends to be small, even as the size of the validation set becomes large. This supported by corollary C.17, since even $1$-nearest neighbor reliably estimates risk in this setting. For the point prediction task, $k^{\star}_{T_{2}}$ grows with $N^{\textup{{val}}}$.
@@ D.3.4 Implementation of Risk Estimation
D.3.5 Monte Carlo Estimation of Test Risk
Following the argument in section D.1, we use the empirical test risk eq. 114 in place of the test risk as ground truth in synthetic experiments. The assumption that $(Z_{m})_{m=1}^{M^{\textup{{test}}}}$ are independent holds by the description of the data generating process, because the $(Y_{m})_{m=1}^{M^{\textup{{test}}}}$ are conditionally independent and $Z_{m}$ is a function of $Y_{m}$. The assumption that $Z_{m}$ is almost surely bounded holds with $\Delta=1$ by our choice of truncated squared loss. Further, in both synthetic experiments, we take $M^{\textup{{test}}}=45000$. Combining these gives that with probability at least $0.95$
@@ D.3.5 Monte Carlo Estimation of Test Risk
fig. 1 shows the absolute difference between each estimator and the empirical test risk across 100 seeds. We see that the difference between the estimators is generally larger than twice eq. 131, and so by the argument in section D.1, we expect our estimate of ground truth to be accurate enough that the difference in performance of the methods is not simply due to error in estimating the ground truth.
D.3.6 Metrics Reported and Additional Figures
Figures 8 and 10 show the relative errors of each estimation, calculated as
@@ D.3.6 Metrics Reported and Additional Figures
D.4 Air Temperature Tasks
We now provide additional details about data source, pre-processing, model fitting and risk estimation for the air temperature dataset. These are identical between the real response experiment and the partially synthetic experiment, except for the bootstrapping procedure described in section D.4.6. We also provide additional experimental results on a grid prediction task for both the bootstrapped and original datasets.
D.4.1 Data Sources
The land surface temperature used is from MODIS Aqua (Wan et al., 2021) and is monthly average land surface temperature on a $0.05$ degree grid. We download monthly average weather station data from the Global Historical Climatology Network (Menne et al., 2018). Latitude and longitude of major United States (US) urban areas are from the 2023 US census gazetteer (United States Census Bureau, 2023). All of these datasets are produced in large part by US government agencies (NASA, NOAA and the Census). While we could not find specific license information, we understand these datasets to be public domain following section 105 of the Copyright Act of 1976.
D.4.2 Data Pre-processing
We assign the land surface temperature at the nearest point (using a spherical approximation to distance between points) to each weather station. The nearest point is found using the scikit-learn implementation of nearest neighbor algorithm using the ‘ball-tree’ (Omohundro, 2009) and ‘Haversine’ metric. Temperatures are converted to Celsius from Kelvin. We remove all rows where the Land Surface temperature is not available. We use the weather station data uploaded to GHCNM as of January 15, 2024. We filter out weather stations outside of the United States (based on the station ID). We also filter out stations with a non-empty quality control flag or no temperature recorded for January 2018 (the month we consider). Finally, we remove stations in Hawaii or Alaska to focus on the continental United States. In total, after this processing, there are 6422 weather stations. We use 70% of the stations for fitting the models, and holdout the remaining 30% estimates. When building the test sites, we remove points outside the United States based on a reverse geocoding lookup with Thampi (2015) to the nearest city. This does not create an exact boundary (since it is based on the nearest city or town and not the country in which the latitude and longitude is based in) but is a good proxy for whether or not a point is in the United States. Figure 11 of the available weather stations for model fitting and validation, colored by monthly average temperature in January 2018.
@@ D.4.3 Loss Function
D.4.4 Model Fitting
Inspired by Hooker et al. (2018), we fit a geographically weighted least squares regression model using the land surface temperature at day and night. In particular, we fit an affine model, with the coefficients, $\beta(S)$ depending on the location that will be predicted at. $\beta(S)$ is selected by solving the weighted least squares problem,
@@ D.4.4 Model Fitting
D.4.5 Risk Estimation Details
The holdout is implemented as in previous experiments (section D.3.4). We estimate the standard error of the holdout empirically by computing the sample standard deviation of the sum of the losses,
@@ D.4.6 Bootstrapping of Residuals
We can then directly estimate the test risk via generating many $Y^{\textup{{test}}}$ at each spatial location (we use 1000 realizations for each city in the 5-metros task) and 1 for each grid point in the grid task (since the error is averaged over test sites this still results in an estimator that is concentrated) in this manner and forming a Monte Carlo estimate as in section D.1.
D.4.7 Estimation of Ground Truth in Bootstrapped Experiment
Following the argument in section D.1, we use the empirical test risk eq. 114 in place of the test risk as ground truth in synthetic experiments. For the assumption that $(Z_{m})_{m=1}^{M^{\textup{{test}}}}$ are independent to hold it is sufficient for $Y^{\textup{{test}}}_{m}$ to be independent, conditioned on the spatial location at which it is observed. This holds based on the data generating process used to construct the synthetic responses: since $Y^{\textup{{test}}}_{m}$ is a noisy observation of the smooth function we fit to the weather stations, plus noise sampled independently from the distribution of residuals. Because the loss is truncated MAE, the $Z_{m}$ are surely bounded by $5$. We use $10000$ samples at each of the 5 test location in estimating the risk. Using these numbers in eq. 117, we arrive at
@@ D.4.7 Estimation of Ground Truth in Bootstrapped Experiment
Following the argument in section D.1, we expect our estimate of ground truth to be good enough to distinguish between the quality of models whose absolute error from the estimated ground truth differs by more than $2\times 0.031=0.062$. Figure 12 shows these absolute errors. We see that for many of the seeds, the difference between the performance of 1NN (orange) and SNN and the holdout is greater than $0.174$. Given that the earlier argument is quite conservative (in the sense that Hoeffding’s inequality is likely to be loose), we therefore can attribute the difference in performance of the methods to indicate that SNN and the holdout are giving better estimates of the ground truth test risk, and the observed difference is not due to error in our estimation of the test risk.
@@
We therefore expect our estimate of ground truth to be very accurate, although we see in fig. 13 that the methods all perform well in estimating the test risk on this task, and so there is likely not a meaningful difference in which approach is used to perform validation.
D.4.8 Results for Grid Prediction with Bootstrapped Data
fig. 13 shows the results for holdout, $1$-nearest neighbor and SNN with the test set each grid point in the map that is located in the continental United States. All 3 methods lead to reasonably accurate estimates of the mean absolute error on this prediction task (within 0.1 degrees of the ground truth error). Based on our theory, we generally expect 1NN and SNN to have small error in grid prediction tasks (at least with sufficient data and the infill assumption being satisfied), while for the holdout it depends on the particular predictive method and distribution of test and validation sites. In this case, it appears for both prediction methods the bias introduced by the use of the holdout is relatively small.
@@ D.4.8 Results for Grid Prediction with Bootstrapped Data
D.4.9 Results for Grid Prediction with Real Data
Table 4 shows the results for holdout, $1$-nearest neighbor and SNN with the test set each grid point in the map that is located in the continental United States. We see good agreement between all three method. This is expected for $1$-nearest neighbor and SNN based on earlier theory (section C.6.5).
@@ D.5 UK Housing Experiment
D.5.1 Data Sources and Pre-processing
We download 2023 price paid data for England and Wales from HM Land Registry (2023). This data is subject to a UK Open Government License (https://www.nationalarchives.gov.uk/doc/open-government-licence/version/3/), which requires citation of the data, but allows both commercial and non-commercial uses. These records contain postal codes for each property sold, the type of property sold, town or city, price paid for the property. We use the type of property variable to filter out all properties that are not flats, and only consider additions (not replacements or deletions) to the dataset and “standard” price paid data (not repossessions, buy-to-lets or other sales labelled as non-standard). Noting that a postal code in the UK corresponds to a very small geographic area, we obtain latitude and longitude data for each sale by looking up the postal code coordinates using the National Statistics Postcode Lookup (Office Of National Statistics, 2024), which we understand to be a product of the UK Census and therefore also subject to an Open Government License. We convert from northing and easting to latitude and longitude using R. We log transform the price variable prior to model fitting as we expect price paid to be non-negative and highly skewed, and so a Gaussian (process) prior would otherwise be almost certainly inappropriate.
D.5.2 Model Fitting
@@
We fit a Gaussian process regression model with prior covariance specified by a sum of two Matérn 3/2 kernels and a zero prior mean on the mean centered log price paid data. We use a sum of Matérn 3/2 kernel in place of the sum of RBF kernel used in Hensman et al. (2013) as we expect there to be places where (log) property prices vary quickly in space, and so the smoothness properties implicitly assumed in using an RBF kernel may be inappropriate. We use 2000 inducing points for the variational approximation. The locations of these points are optimized jointly with model parameters when maximizing the evidence lower bound. We use the closed form for the optimal variational posterior (given a set of inducing points) provided in Titsias (2009), and perform maximization of the evidence lower bound using L-BFGS.
@@
The locations of the inducing points are initialized by the greedy procedure suggested in Burt et al. (2020), which is essentially equivalent to a partially pivoted Cholesky decomposition recommended earlier in the Gaussian process approximation literature (Foster et al., 2009). The initial prior variance of both kernels is set to be equal to the variance in the training data; the lengthscales of one kernel (intended to model regional price trends) are initialized to twice the standard deviation in the location data (in radians), while the scale of the other kernel (intended to model local trends) is initialized to half the standard deviation in the location data. The likelihood standard deviation is initialized to be 0.1$\times$ the standard deviation in the log price paid in the training data.
D.5.3 Estimation of Ground Truth
We next describe why we might expect empirical test risk to provide a reasonable estimate of ground truth in this problem, and in particular we discuss the assumptions that justify the use of empirical test risk in relation to section D.1. We consider a truncated loss $\ell(a,b)=\min(0,10^{6})$, and so $Z_{m}$ is almost surely bounded by $10^{6}$. The assumption that the $(Z_{m})_{m=1}^{M^{\textup{{test}}}}$ are independent would be implied by independence of $(Y_{m})_{m=1}^{M^{\textup{{test}}}}$ that is: that given the location at which a flat is sold, any remaining randomness in the observed sale prices is independent. Concretely, we might think of the randomness in sales price, $\epsilon^{\text{test}}_{m}$ in our model, as coming from aspects of the sale process of the house, such as who happens to see the advertisement for a house, and we assume these are independent for each house when constructing our estimate of the ground truth.
If the conditional independence assumption proposed above holds, then following section D.1 we have that
@@ D.5.3 Estimation of Ground Truth
D.5.4 Risk Estimation Details
Holdout, 1NN and SNN are run as in the air temperature experiments (section D.4.5). In particular, we use a failure probability of $\delta=0.1$ for SNN and nearest neighbor calculations are done with respect to Haversine distance to account for Earth’s curvature. We use a fixed $\delta$ since we only discuss a method applicable to estimating fill distance on the unit cube, not on a subset of the sphere. $\Delta=\pounds 1,000,000$ is used in selecting the number of neighbors as this is an upper bound on the truncated loss. We use a Lipschitz constant of £1,000/km as £1/km seems implausibly small.
D.6 Wind Speed Prediction Experiment
@@
D.6.1 Data Sources and Pre-processing
We download daily wind speed readings from weather stations from the Global Historical Climate Network (Menne et al., 2012). As this dataset was constructed by NOAA employees, we understand it to be public domain following section 105 of the Copyright Act of 1976. We filter out weather stations outside the continental US, as well as any weather stations that do not contain daily average wind speed readings. We look only at wind speed data from January in the prediction task, and years 2000–2024. There is a weather station at Chicago O’Hare which we remove from the training and validation data and use as the test set.
For each replicate used to form fig. 13 we split off a training set containing 80% of weather stations, and a validation set containing the remaining 20%. The number of observations in the training and validation set varies (because different weather stations may be online for a different number of days in January in previous years), but this leads to on the order of 580000 training observation and 126000 validation observations. Each training and validation point is a triple containing latitude, longitude and average wind speed. We perturb the latitude and longitude (in degrees) of validation points by a Gaussian random variable with standard deviation $10^{-12}$, which is essentially equivalent to using random tie-breaking in the nearest neighbor algorithms. We expect this has a significant impact on 1NN compared to the version discussed in the paper, because there may be many observations from the nearest weather station to Chicago O’Hare. While this would unlikely be done in practice, random tie-breaking, or tie-breaking by selecting the first nearest neighbor according to some other ordering are common and would lead to similar outcomes as the results presented here (but higher variance than averaging over all neighbors that are equally close). The latitude and longitude are then converted to radians for the analysis.
D.6.2 Loss Function